OpenStudy (anonymous):

WHO WANTS A MEDAL What is the equation for a line that passes through the points (-2,2) and (4,10) y= -2x + 6 y= 2x - 6 y= -2x + 2 y= 2x + 2

OpenStudy (loser66):

do you want the medal?

OpenStudy (loser66):

@PixieDust1 this is your job

OpenStudy (anonymous):

No. I give one for answer.

OpenStudy (anonymous):

dentify the slope of the line below. 7x + 4y = 16

OpenStudy (pixiedust1):

Lol uh oh! =P Oh, so here goes so y=mx+b That Loser kindly devoted his time to teaching me :D

OpenStudy (anonymous):

4 -4 7/4 -7/4

OpenStudy (pixiedust1):

So, The formula is y=mx+b

OpenStudy (pixiedust1):

And then, this is when Loser freaks out because I forget what to do next =P

OpenStudy (pixiedust1):

Oh wait! I know how to do the second one!

OpenStudy (loser66):

@PixieDust1 take advantage on this problem, the asker just wants the answer and this fact give you a good opportunity to practice without any cost. If you are right, good for you, If you are wrong, no cost. Ha!! why not taking it, right?

OpenStudy (pixiedust1):

Okay!!! I'll certainly try! But you better check mah work!

OpenStudy (anonymous):

i just need help with second one now, please

OpenStudy (pixiedust1):

So, we wanna isolate the y. Therefore, let's get the 7x away from it 7x+4y=16 -7x -7x Now, its 4y=-7x+16 Am I right so far?

OpenStudy (loser66):

yup

OpenStudy (pixiedust1):

Yay! Okay, so King, what do you think the next step would be?

OpenStudy (pixiedust1):

Remember, we wanna isolate the y.

OpenStudy (loser66):

@PixieDust1 good guy and an A student, proud of you,!!

OpenStudy (pixiedust1):

Yay!! :D I'm proud of myself too lol!

OpenStudy (pixiedust1):

I wouldn't be able to do this stuff if it wasn't for all that time Loser spent tutoring me!